OpenStudy (anonymous):

solve using the substitution method x+y=6 and y=x-4

OpenStudy (anonymous):

i have gave a attachment with your previous question similarly u will find here solution will be y=2,x=4

OpenStudy (anonymous):

if y = x-4, then x+y = 6 becomes x+x-4 =6 2x-4 =6 2x =10 x =5

OpenStudy (anonymous):

x-2y=5 and 2x-y=16 solve using the substitution methd

OpenStudy (anonymous):

chloe, the answer is x =5, y =1 and not y=2,x=4 as dipankarstudy said

OpenStudy (anonymous):

x=5+2y putting it in 2nd equation 10+4y-y=16 3y=6 y=2 and x=9

OpenStudy (anonymous):

yes i made a mistake...

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

thanks dhatra...